OpenStudy (livya15):

I need help with understanding logarithms, please. The directions: Find the Value of x. ln(2x-1)+2=0

OpenStudy (solomonzelman):

\[\ln(2x-1)+2=0\]\[\ln(2x-1)=-2\]\[\ln(2x-1)=-2\times \ln(e)\]\[\ln(2x-1)= \ln(e^{-2})\]\[2x-1=e^{-2}\]\[2x-1=\frac{1}{e^2}\]\[2x=\frac{1}{e^2}+1\]\[2x=\frac{1}{e^2}+\frac{e^2}{e^2}\]\[2x=\frac{e^2+1}{e^2}\]\[x=\frac{e^2+1}{2e^2}\]

OpenStudy (solomonzelman):

you can use 2.7 approximation for e, or leave it the way it is now.

OpenStudy (livya15):

Thankyou SOOO much....that makes a lot of sense :)

OpenStudy (solomonzelman):

You welcome:) Glad to help !

OpenStudy (solomonzelman):

Fixed your color :)